1. How much energy is needed to change the temperature of 50.0 g of water by 15.0oC? (Specific heat capacity of water = 4.18 J/g °C)


  






2. How many grams of water can be heated from 20.0 oC to 75oC using 12500.0 Joules? (Specific heat capacity of water = 4.18 J/g °C) 







 3. What is the total temperature change after 840 Joules is absorbed by 10.0g of water? (Specific heat capacity of water = 4.18 J/g °C)




4. How much energy is needed to raise the temperature of a 850 g block of  aluminum from 22.8oC to 94.6oC? (Specific heat capacity of aluminum is 0.900 J/goC)









5. You are given the same amount of iron and lead (5g of each). If you want to heat them both up by 20oC, which substance would have a higher amount of heat transferred? Explain WHY. (HINT: Refer to the table) 









6. You are given 50 g of lead and 100 g of lead. Which sample will require MORE heat energy to raise the temperature by 10oC? Why?

















1
Expert's answer
2021-04-07T11:48:04-0400

1)

"Q= m\u00d7c\u00d7dT"

= 50×4.18× 15= 3135J

2)m= 12500/(4.18× 55)

m=54.4g

3)dT=840/(10×4.18)

dT=200C

4)Q=840×4.18×71.8

= 252104.16J

5)Iron would have a higher amount of heat because it has a higher specific heat capacity

6)the 100g mass will require more energy to raise it's temperature by 100C